# TopER: Topological Embeddings in Graph Representation Learning

### v2

@@ 4.1 Computational Complexity
Thus, the overall complexity of TopER predominantly hinges on the graph filtration process, summing up to $O(n\times(|\mathcal{V}|+|\mathcal{E}|))$ where $|\mathcal{V}|\gg n$. As we will show in the next section, the runtime costs of TopER are notably low, demonstrating its practicality and efficiency for large-scale applications.
@@ 4.2 Stability Results
This section states our theorems on the stability of TopER. In the following, $\mathcal{W}_{p}(.,.)$ represents $p$-Wasserstein distance, and $\rm{PD}_{k}(\mathcal{X},f)$ represents $k^{th}$ persistence diagram of $\mathcal{X}$ with sublevel filtration with respect to $f$. Similarly, .$\|.\|_{p}$ represents $L^{p}$-norm and $d_{p}(.,.)$ represents $l_{p}$-distance in $\mathbb{R}^{m}$. We fix a threshold set =$\mathcal{I}=\{\epsilon_{i}\}_{i=1}^{n}$ for both functions to keep the exposition simple. Further, to keep the setting general, we use the pairs $\{(\beta_{0}(\epsilon_{i}),\beta_{1}(\epsilon_{i}))\}_{i=1}^{n}$ in $\mathbb{R}^{2}$ to fit the least squares line $y=a+bx$ defining $\rm{TE}(\mathcal{X})=(a,b)$.
Theorem 1.
@@
By adapting the above results to the graph setting, when two metric graphs $\mathcal{G}_{1},\mathcal{G}_{2}$ are close in Gromov-Hausdorff sense, one can obtain a similar stability result for the filtrations of $\mathcal{G}_{i}$ induced by the same filtration function. Due to space limitations, details and the proof of the theorem are given in Appendix C.
@@ 5 Experiments
We evaluate the performance of TopER in classification, clustering and visualization. Our Python implementation is available at https://anonymous.4open.science/r/TopER-AA38.
@@ 5.1 Experimental Setup
Datasets. We conduct experiments on nine benchmark datasets for graph classification. These are (i) the molecule graphs of BZR, and COX2 (Mahé and Vert,, 2009); (ii) the biological graphs of MUTAG and PROTEINS (Kriege et al.,, 2012); and (iii) the social graphs of IMDB-Binary (IMDB-B), IMDB-Multi (IMDB-M), REDDIT-Binary (REDDIT-B), and REDDIT-Multi-5K (REDDIT-5K) (Yanardag and Vishwanathan,, 2015). Finally, the OGBG-MOLHIV is a large molecular property prediction dataset, part of open graph benchmark (OGB) datasets (Hu et al.,, 2020). Details are given in Table 1.
@@
Model Setup and Metrics. We employ a rigorous experimental setup to ensure a fair comparison and the selection of the best graph classification model. We begin by applying BatchNormalization to the input features to maintain consistent scaling. We employ a 90/10 train-test split, adopt the StratifiedkFold strategy, and present the average accuracy from ten-fold cross-validation across all our models. We employ accuracy as the evaluation metric, a widely utilized performance measure within graph classification tasks (Errica et al.,, 2020).
Filtration functions. In TopER, we use both node and edge filtrations during filtration (Definition 1). Alongside with popularity, we apply degree, closeness, and degree centrality  (Evans and Chen,, 2022) as node filtration functions and Forman- and Ollivier-Ricci functions (Lin et al.,, 2011) as edge filtration functions. We also use atomic weight as a node function for molecular and biological datasets (BZR, COX2, and MUTAG), and node attributes (PROTEINS). We utilized the t-test to assess the statistical significance of each function and applied the Lasso method for regularization, setting the cross-validation parameter to $cv=10$. Functions were retained in the model only if they achieved p-values less than 0.05 in the t-test and had non-zero coefficients in the Lasso model (James et al.,, 2023). This approach ensures that the selected filtration functions contribute statistically significant and regularized features to the model. Incorporating additional filtration functions can enhance TopER’s ability to analyze graphs from diverse perspectives. However, as we will next illustrate in Table 6, TopER demonstrates strong performance even in its most basic form using the simple and scalable node degree function. This balance of performance and simplicity suits our scalability philosophy; we avoid complex and costly schemes for learning dataset-specific activation functions and homogenize the filtration step in all datasets.
Classifier. We utilize a Multilayer Perceptron (MLP) in our graph classification task. The hyperparameters are detailed in Section A.1.
Figure 4: TopER run time for synthetic power law graphs (Holme and Kim,, 2002) with node degree filtration. The mean node degree is $30$, and 100 filtration steps are used.
Baselines. We compare our method with 19 state-of-the-art and recent models in graph classification, including variants of graph neural networks: six GNNs including GCN, DGCNN, Diffpool, ECC, GIN, GraphSAGE which are compared in (Errica et al.,, 2020) (best results of these six GNNs are given in the 6 GNNs row), FF-GCN (Paliotta et al.,, 2023); topological methods: DMP (Bodnar et al.,, 2021), FC-V (O’Bray et al.,, 2021), WWLS (Fang et al.,, 2023), MP-HSM (Loiseaux et al.,, 2024) and EMP (Chen et al., 2024a, ); GNNs enhanced with data augmentation methods: SubMix (Yoo et al.,, 2022), G-Mix (Han et al.,, 2022), and EPIC (Heo et al.,, 2024); GNNs enhanced with contrastive learning methods: RGCL (Li et al.,, 2022), AutoGCL (Yin et al.,, 2022), TopoGCL (Chen et al., 2024b, ) and prototype-based methods: PGOT (Qian et al.,, 2024).
Table 2: Graph Classification. Accuracy results on eight benchmark datasets using 10-fold CV. Baseline results are sourced from the corresponding papers. The best performance is highlighted in bold blue, while the second-best performance is underlined. The final column presents the average deviation of each model’s performance from the best result across all datasets.
Model BZR COX2 MUTAG PROTEINS IMDB-B IMDB-M REDDIT-B REDDIT-5K Avg.$\downarrow$
6 GNNs (Errica et al.,, 2020) 80.42$\pm$2.07 75.80$\pm$3.70 71.20$\pm$3.90 49.10$\pm$3.50 89.90$\pm$1.90 56.10$\pm$1.60 4.13
DMP (Bodnar et al.,, 2021) 84.00$\pm$8.60 75.30$\pm$3.30 73.80$\pm$4.50 50.90$\pm$2.50 86.20$\pm$6.80 51.90$\pm$2.10 4.20
FC-V (O’Bray et al.,, 2021) 85.61$\pm$0.59 81.01$\pm$0.88 87.31$\pm$0.66 74.54$\pm$0.48 73.84$\pm$0.36 46.80$\pm$0.37 89.41$\pm$0.24 52.36$\pm$0.37 4.01
SubMix (Yoo et al.,, 2022) 86.34$\pm$2.00 84.68$\pm$3.70 80.99$\pm$0.60 67.80$\pm$2.00 70.30$\pm$1.40 46.47$\pm$2.50 6.15
G-Mix (Han et al.,, 2022) 84.15$\pm$2.30 83.83$\pm$2.10 81.96$\pm$0.60 66.28$\pm$1.10 69.40$\pm$1.10 46.40$\pm$2.70 6.91
RGCL (Li et al.,, 2022) 84.54$\pm$1.67 79.31$\pm$0.68 87.66$\pm$1.01 75.03$\pm$0.43 71.85$\pm$0.84 49.31$\pm$0.42 90.34$\pm$0.58 56.38$\pm$0.40 3.56
AutoGCL (Yin et al.,, 2022) 86.27$\pm$0.71 79.31$\pm$0.70 88.64$\pm$1.08 75.80$\pm$0.36 72.32$\pm$0.93 50.60$\pm$0.80 88.58$\pm$1.49 56.75$\pm$0.18 3.08
FF-GCN (Paliotta et al.,, 2023) 89.00$\pm$5.00 78.00$\pm$8.00 71.00$\pm$4.00 62.00$\pm$1.00 63.00$\pm$8.00 11.53
WWLS (Fang et al.,, 2023) 88.02$\pm$0.61 81.58$\pm$0.91 88.30$\pm$1.23 75.35$\pm$0.74 75.08$\pm$0.31 51.61$\pm$0.62 2.26
EPIC (Heo et al.,, 2024) 88.78$\pm$2.30 85.53$\pm$1.60 82.44$\pm$0.70 69.06$\pm$1.00 71.70$\pm$1.00 47.93$\pm$1.30 4.67
EMP (Chen et al., 2024a, ) 88.79$\pm$0.63 72.78$\pm$0.54 74.44$\pm$0.45 48.01$\pm$0.42 91.03$\pm$0.22 54.41$\pm$0.32 2.97
MP-HSM (Loiseaux et al.,, 2024) 77.10$\pm$3.00 85.60$\pm$5.30 74.60$\pm$2.10 74.80$\pm$2.50 47.90$\pm$3.20 4.67
TopoGCL (Chen et al., 2024b, ) 87.17$\pm$0.83 81.45$\pm$0.55 90.09$\pm$0.93 77.30$\pm$0.89 74.67$\pm$0.32 52.81$\pm$0.31 90.40$\pm$0.53 1.76
PGOT (Qian et al.,, 2024) 87.32$\pm$3.90 82.98$\pm$5.21 92.63$\pm$2.58 73.21$\pm$2.59 62.90$\pm$3.05 51.33$\pm$1.76 3.85
TopER 90.13$\pm$4.14 82.01$\pm$4.59 90.99$\pm$6.64 74.58$\pm$3.92 73.20$\pm$3.43 50.00$\pm$4.02 92.70$\pm$2.38 56.51$\pm$2.22 1.60
@@
Runtime. TopER is highly scalable and can be applied to a 100K node graph in  2 minutes (see Figure 4). Our small network experiments took about two days in a shared resource setting, whereas the OGBG-MOLHIV experiments took 16.5 hours. One of the most demanding datasets, REDDIT-5K, requires $34.6$ hours to calculate all node and edge functions. The runtime of our methods is dominated by the computation of node functions such as closeness and Riccis. Using approximate values for centrality metrics instead could greatly decrease computation time (Brandes and Pich,, 2007). Since this is not our current focus, we leave it as future work.
@@ 5.2 Graph Classification Results
Table 2 shows the accuracy results for the given models. We use the reported results in the corresponding references for each model. “--" entries in the table mean the reference did not report any result for that dataset. In (Errica et al.,, 2020), the authors compare the six most common GNNs on the graph classification task (see the GNNs row). The last column summarizes each model’s overall performance. We report the average of the differences between each model’s performance and the best performance in the column across all datasets. If a model’s performance is missing for a dataset, it is excluded from the average computation for the model.
@@
GIN-VN (Xu et al.,, 2018) 77.80$\pm$1.82
HGK-WL (Togninalli et al.,, 2019) 79.05$\pm$1.30
WWL (Borgwardt et al.,, 2020) 75.58$\pm$1.40
PNA (Corso et al.,, 2020) 79.05$\pm$1.32
DGN (Beaini et al.,, 2021) 79.70$\pm$0.97
GraphSNN (Wijesinghe and Wang,, 2021) 79.72$\pm$1.83
GCN-GNorm (Cai et al.,, 2021) 78.83$\pm$1.00
Graphormer (Ying et al.,, 2021) 80.51$\pm$0.53
Cy2C-GCN (Choi et al.,, 2022) 78.02$\pm$0.60
GAWL (Nikolentzos and Vazirgiannis,, 2023) 78.34$\pm$0.39
LLM-GIN (Zhong et al.,, 2024) 79.22$\pm$±NA
GMoE-GIN (Wang et al.,, 2024) 76.90$\pm$0.90
TopER 80.21$\pm$0.15
OGBG-MOLHIV results. To evaluate our model’s performance on large datasets, we compare it with recently published models on OGBG-MOLHIV dataset, as shown in Table 3. The performances of these models are listed in chronological order based on their publication dates, with baseline performances reported from (Choi et al.,, 2022; Ying et al.,, 2021) or the respective model’s references. In Section A.2, we give further details for TopER performance and contribution of each function on this dataset. TopER achieves the second-best result on the MOLHIV dataset, while the top-performing model requires learning a significantly larger model with 119.5 million parameters.
@@
PH 88.4$\pm$0.6 82.0$\pm$0.6 74.0$\pm$0.4 69.5$\pm$0.5 46.5$\pm$0.3 54.1$\pm$0.1
TopER 90.1$\pm$4.1 82.0$\pm$4.6 74.6$\pm$3.9 73.2$\pm$3.4 50.0$\pm$4.0 56.5$\pm$2.2
TopER vs. PH. TopER-based models are more accurate than those based on Persistent Homology. We present a comparison across six datasets in Table 4. The PH results are obtained from (Cai,, 2021), which extensively examines PH methods in graph classification tasks using four common filtration functions (centrality, degree, fiedler_s, and Ricci), coupled with four popular vectorization techniques. For each dataset, these methods yield 16 outcomes, and we bring in the top-performing combination of filtration and vectorization for each dataset. The results show that TopER outperforms PH in all benchmark datasets.
@@ 5.3 Graph Clustering Results
We employ cluster quality metrics to assess the embeddings of graphs sourced from all datasets in Table 2. The embeddings are labeled with their respective dataset memberships, and we assume that good embeddings will have graphs of the same dataset clustered together. We evaluate embeddings based on three widely used clustering metrics: Silhouette (SILH), Calinski-Harabasz (CH), and Davies-Bouldin (DB) (Gagolewski et al.,, 2021). Table 5 compares the clustering performance of TopER and Spectral Zoo (Jin and Zafarani,, 2020), which is, to our knowledge, the only model that allows low-dimensional graph embeddings. Detailed results are provided in Section A.3. The findings demonstrate that the embeddings generated by TopER outperform those created by Spectral Zoo. This is evident from the superior cluster quality metrics observed for five out of eight datasets in the case of Silhouette and CH, and for all eight datasets in the case of DB.
Table 5: Clustering Performances. Comparison of Spectral Zoo vs. TopER. The detailed results are given in Section A.3.
@@ 5.4 Graph Visualization
In the case of a single filtration function, TopER creates $2D$ graph embeddings $(a,b)$ that can be visualized with ease (see Figure 1). Traditional dimensionality reduction techniques such as PCA can be used to visualize point cloud data, but accurately depicting graph data has historically been a significant challenge (Giovannangeli et al.,, 2020). To our knowledge, the only model that allows graph visualization is the GraphZoo (Jin and Zafarani,, 2020).
TopER creates highly interpretable graph visualizations. To recall, the pair $(a,b)$ represents the coefficients of the best-fitting function $L(x)=a+bx$, where $a$ is the pivot (y-intercept) and $b$ is growth (the slope). Specifically, the pivot $a$ reflects graph connectivity, while $b$ reflects the growth rate of edges/nodes for the filtration function. In particular, a higher value of $a$ corresponds to a more interconnected graph. As we demonstrate in Figure 7, graph connectivity and community structure can be analyzed using three types of pivot behavior. In the following, we illustrate how these quantities can be employed to interpret our two-dimensional representations of the graph datasets.
@@
Figure 5: TopER visualizations of the PROTEINS dataset with Ollivier-Ricci edge filtration, and the BZR dataset with degree centrality node filtration. Each point corresponds to an individual graph.
@@
In Figure 1(b) of the MUTAG dataset, class B has a higher growth rate and smaller pivot than the red class. This shows that the class is growing faster than class A with respect to the closeness function in the MUTAG dataset, i.e., the graph has a low diameter. Similarly, in contrast, in the PROTEINS dataset (Figure 5(a)), the growth rates are similar for both classes ($\sim 1.5-1.7$), but the pivot (initial graph size) is smaller in class A. This implies that class A has fewer edges in relation to the number of nodes. Such patterns, as described in Section B.4, can reveal key insights into graph topology. In a similar vein, TopER visualizations can be used for anomaly detection. For example, in Figure 1(a), an outlier PROTEINS graph alone has a positive pivot in the figure and appears as the rightmost data point.
@@
More importantly, TopER homogenizes graph embeddings, allowing us to compare graphs across datasets, which may open new paths in training graph foundation models. For example, Figure 1(a) visualizes graphs of three datasets on the same panel where we see that Mutag and COX2 differ in their pivot only. The similarity is not surprising; MUTAG and COX2 are datasets of molecular graphs where nodes are atoms and edges are chemical bonds. As the molecules in both datasets have similar types of atoms and bond configurations (e.g., ring structures), TopER captures these similarities, leading to similar embeddings.
5.5 Ablation Study
In an ablation study, we evaluate the individual performance of each function as well as their combined effect on classification. As shown in Table 6, the common filtration functions we employ from TDA exhibit strong individual performance. Moreover, when combined, they synergistically enhance overall performance.
BZR 82.22$\pm$2.13 82.20$\pm$3.42 81.48$\pm$1.99 82.73$\pm$2.12 80.75$\pm$1.73 80.99$\pm$1.48 82.23$\pm$2.12 90.13$\pm$4.14
COX2 75.38$\pm$3.96 69.21$\pm$8.19 67.90$\pm$7.96 73.88$\pm$5.02 70.46$\pm$7.28 73.03$\pm$4.21 69.82$\pm$8.27 82.01$\pm$4.59
MUTAG 76.61$\pm$7.87 77.66$\pm$6.12 80.88$\pm$4.79 74.97$\pm$6.40 80.85$\pm$9.25 82.46$\pm$7.84 73.45$\pm$8.01 90.99$\pm$6.64
PROTEINS 67.66$\pm$3.16 70.71$\pm$4.41 69.01$\pm$4.24 69.01$\pm$3.48 72.96$\pm$3.47 71.25$\pm$2.66 73.59$\pm$3.33 74.58$\pm$3.92
IMDB-B 73.00$\pm$4.49 71.90$\pm$3.48 72.60$\pm$4.20 73.10$\pm$4.18 69.80$\pm$2.44 66.40$\pm$3.35 - 73.20$\pm$3.43
IMDB-M 48.47$\pm$3.90 47.87$\pm$3.07 48.33$\pm$3.49 47.93$\pm$2.88 48.13$\pm$4.11 43.60$\pm$3.17 - 50.00$\pm$4.02
REDDIT-B 76.70$\pm$3.69 79.35$\pm$3.46 78.10$\pm$3.23 79.55$\pm$2.20 72.35$\pm$2.91 68.20$\pm$2.28 - 92.70$\pm$2.38
REDDIT-5K 42.85$\pm$1.74 50.87$\pm$2.63 50.03$\pm$1.49 47.01$\pm$1.89 50.27$\pm$1.92 45.81$\pm$2.08 - 56.51$\pm$2.22

### v3

@@ 3.1 Computational Complexity
Thus, the overall complexity of TopER predominantly hinges on the graph filtration process, summing up to $O(n\times(|\mathcal{V}|+|\mathcal{E}|))$ where $|\mathcal{V}|\gg n$. As we will show in the next section, the runtime costs of TopER are notably low, making it practical and efficient for large-scale applications.
@@ 3.2 Stability Results
This section states our theorems on the stability of TopER. In the following, $\mathcal{W}_{p}(.,.)$ represents $p$-Wasserstein distance, and $\rm{PD}_{k}(\mathcal{X},f)$ represents $k^{th}$ persistence diagram of $\mathcal{X}$ with sublevel filtration with respect to $f$. Similarly, $\|.\|_{p}$ represents $L^{p}$-norm and $d_{p}(.,.)$ represents $l_{p}$-distance in $\mathbb{R}^{m}$. We fix a threshold set $\mathcal{I}=\{\epsilon_{i}\}_{i=1}^{n}$ for both functions to keep the exposition simple. Further, to keep the setting general, we use the pairs $\{(\beta_{0}(\epsilon_{i}),\beta_{1}(\epsilon_{i}))\}_{i=1}^{n}$ in $\mathbb{R}^{2}$ to fit the least squares line $y=a+bx$ defining $\rm{TE}(\mathcal{X})=(a,b)$.
Theorem 3.4.
@@
The following lemmas are essential for proving the theorem and corollary above.
Lemma 3.6.
(ST, 20) Let $\mathcal{X}$ be a compact metric space, and $f,g:\mathcal{X}\to\mathbb{R}$ be two filtration functions. Then, for any $p\geq 1$, we have $\mathcal{W}_{p}(\rm{PD}_{k}(\mathcal{X},f),\rm{PD}_{k}(\mathcal{X},g))\leq\|f-g\|_{p}$
The next lemma is on the stability of Betti curves by (DG, 23) [Proposition 1].
Lemma 3.7.
(DG, 23) Let $\beta_{k}(\mathcal{X})$ is the $k^{th}$ Betti function obtained from the persistence module $\rm{PM}_{k}(\mathcal{X})$.
By adapting the above results to the graph setting, when two metric graphs $\mathcal{G}_{1},\mathcal{G}_{2}$ are close in the Gromov-Hausdorff sense, one can obtain a similar stability result for the filtrations of $\mathcal{G}_{i}$ induced by the same filtration function. Due to space limitations, details and the proofs are given in Appendix˜C.
@@ 4 Experiments
We evaluate the performance of TopER in classification, clustering and visualization. Our Python implementation is available at https://github.com/AstritTola/TopER.
@@ 4.1 Experimental Setup
Datasets. We conduct experiments on nine benchmark datasets for graph classification. These
@@
are (i) the molecule graphs of BZR, and COX2 (MV, 09); (ii) the biological graphs of MUTAG and PROTEINS (KM, 12); and (iii) the social graphs of IMDB-Binary (IMDB-B), IMDB-Multi (IMDB-M), REDDIT-Binary (REDDIT-B), and REDDIT-Multi-5K (REDDIT-5K) (YV, 15). Finally, the OGBG-MOLHIV is a large molecular property prediction dataset, part of the open graph benchmark (OGB) datasets (HFZ+, 20). Data statistics are given in Table˜1.
@@
Runtime. TopER is highly scalable and can be applied to a 100K node graph in less than 2 minutes (see Figure˜5). Our small network experiments took about two days in a shared resource setting, whereas the OGBG-MOLHIV experiments took 7.85 hours. One of the most demanding datasets, REDDIT-5K, requires $2.91$ hours to calculate all node and edge functions. The runtime of our methods is dominated by the computation of node functions such as closeness and Riccis blue(see Section˜A.6). Using approximate values for centrality metrics instead could greatly decrease computation time (BP, 07). Since this is not our current focus, we leave it as future work.
Model Setup and Metrics. We employ a rigorous experimental setup to ensure a fair comparison and the selection of the best graph classification model. We begin by applying BatchNormalization to the input features to maintain consistent scaling. We employ a 90/10 train-test split, adopt the StratifiedkFold strategy, and present the average accuracy from ten-fold cross-validation across all our models. We employ accuracy as the evaluation metric, a widely utilized performance measure within graph classification tasks (EPBM, 20).
Filtration functions. In TopER, we use both node and edge filtrations (Definition˜3.1). Alongside popularity, we apply degree, closeness, and degree centrality (EC, 22) as node filtration functions and Forman- and Ollivier-Ricci functions (LLY, 11) as edge filtration functions. We also use atomic weight as a node function for molecular and biological datasets (BZR, COX2, and MUTAG), and node attributes (PROTEINS). We utilized the t-test to assess the statistical significance of each function and applied the Lasso method for regularization. Functions were retained in the model only if they achieved p-values less than 0.05 in the t-test and had non-zero coefficients in the Lasso model (JWH+, 23). This approach ensures that the selected filtration functions contribute statistically significant and regularized features to the model. Incorporating additional filtration functions can enhance TopER’s ability to analyze graphs from diverse perspectives. However, as we will next illustrate in Table 5, TopER demonstrates strong performance even in its most basic form using the simple and scalable node degree function. This balance of performance and simplicity suits our scalability philosophy; we avoid complex and costly schemes for learning dataset-specific activation functions and homogenize the filtration step in all datasets.
Classifier. We utilize a Multilayer Perceptron (MLP) in our graph classification task. The hyperparameters are detailed in Section˜A.7.
@@ 4.2 Graph Classification Results
Table 2: Graph Classification. Accuracy results on eight benchmark datasets. Best results are in bold blue, second-best are underlined. The final column shows each model’s average deviation from the best per dataset.
[…]
Baselines. We compare our method with 22 state-of-the-art and recent models in graph classification, including variants of graph neural networks: six GNNs including GCN, DGCNN, Diffpool, ECC, GIN, GraphSAGE which are compared in (EPBM, 20) (best results of these six GNNs are given in the 6 GNNs row), FF-GCN (PAMF, 23); topological methods: DMP (BCL, 21), FC-V (ORB, 21), WWLS (FHSK, 23), MP-HSM (LSC+, 23) and EMP (CSA+, 23); GNNs enhanced with data augmentation methods: SubMix (YSK, 22), G-Mix (HJLH, 22), and EPIC (HLAK, 24); GNNs enhanced with contrastive learning methods: RGCL (LWZ+, 22), AutoGCL (YWH+, 22), TopoGCL (CFG, 24); prototype-based methods: PGOT (QTLL, 24); pooling methods: Top-k  (GJ, 19), SAGPool  (LLK, 19), DiffPool  (YYM+, 18), MinCutPool  (BGA, 19) and structural encoder: GPSE (CLL+, 24).
Table˜2 shows the accuracy results for the given models. We use the reported results in the corresponding references for each model. “-" entries in the table mean the reference did not report any result for that dataset. In (EPBM, 20), the authors compare the six most common GNNs on the graph classification task (see the GNNs row). The last column summarizes each model’s overall performance. We report the average of the differences between each model’s performance and the best performance in the column across all datasets. If a model’s performance is missing for a dataset, it is excluded from the average computation for the model.
@@
[…]
OGBG-MOLHIV results. To evaluate our model’s performance on large datasets, we compare it with recently published models on the OGBG-MOLHIV dataset, as shown in Table˜3. The performances of these models are listed in chronological order based on their publication dates, with baseline performances reported from (CPWC, 23; YCL+, 21) or the respective model’s references. In Section˜A.1, we give further details for TopER performance and the contribution of each function on this dataset. TopER achieves the second-best result on the MOLHIV dataset, while the top-performing model requires learning a significantly larger model with 119.5 million parameters.
TopER vs. PH. TopER consistently outperforms Persistent Homology methods in both accuracy and computational efficiency. As shown in Table˜4, we compare against the best PH results reported in (Cai, 21), which evaluates 16 combinations of four filtration functions and four vectorization techniques per dataset. TopER achieves higher accuracy on all six benchmarks. In terms of runtime, TopER is over 10 times faster than PH on large graphs like Reddit-5K, while maintaining high performance (see Table˜8 and Section˜A.2).
@@
[…]
Ablation Studies.
We have conducted three ablation studies. In the first one, we evaluated the individual performance of each function as well as their combined effect on classification. As shown in Table˜5, the common filtration functions we employ from TDA exhibit strong individual performance. Moreover, when combined, they synergistically enhance overall performance. This is not surprising, as different filtration functions, such as atomic weight or Ricci curvature, generate distinct hierarchies and node-edge distributions, resulting in diverse connectivity patterns throughout the filtration sequence. This diversity is analogous to viewing an object from multiple angles. Hence, integrating these complementary perspectives improves performance by offering a richer and more varied representation of the graph structure, allowing the model to capture more intricate features. The other two ablation studies are provided in the Section˜A.4, which examines the effect of the number of thresholds on TopER’s performance, while Section˜A.5 analyzes the impact of the number of filtration functions used.
[…]
@@ 4.3 Graph Clustering Results
We employ cluster quality metrics to assess the embeddings of graphs sourced from all datasets in Table 2. The embeddings are labeled with their respective dataset memberships, and we assume that good embeddings will have graphs of the same dataset clustered together. We evaluate embeddings based on three widely used clustering metrics: Silhouette (SILH), Calinski-Harabasz (CH), and Davies-Bouldin (DB) (GBC, 21). Table˜6 compares the clustering performance of TopER and Spectral Zoo (JZ, 20), which is, to our knowledge, the only model that allows low-dimensional graph embeddings. Detailed results are provided in Section˜A.3. The findings demonstrate that the embeddings generated by TopER outperform those created by Spectral Zoo. This is evident from the superior cluster quality metrics observed for five out of eight datasets in the case of Silhouette and CH, and for all eight datasets in the case of DB.
Table 6: Clustering Performances. Comparison of Spectral Zoo vs. TopER. The detailed results are given in Section˜A.3.
@@ 4.4 Graph Visualization
In the case of a single filtration function, TopER creates $2D$ graph embeddings $(a,b)$ that can be visualized with ease (see Figure˜1). Traditional dimensionality reduction techniques
(a) PROTEINS
(b) BZR
@@
Figure 4: TopER visualizations of the PROTEINS dataset with O.Ricci edge filtration, and the BZR dataset with degree centrality node filtration. Each point corresponds to an individual graph.
such as PCA can be used to visualize point cloud data, but accurately depicting graph data has historically been a significant challenge (GBGA, 20). To our knowledge, until TopER, the only model that allowed graph visualization was the GraphZoo (JZ, 20).
TopER creates highly interpretable graph visualizations. To recall, the pair $(a,b)$ represents the coefficients of the best-fitting function $L(x)=a+bx$, where $a$ is the pivot (y-intercept) and $b$ is the growth (the slope). Specifically, the pivot $a$ reflects graph connectivity, while $b$ reflects the growth rate of edges/nodes for the filtration function. In particular, a higher value of $a$ corresponds to a more interconnected graph. As we demonstrate in Appendix Figure˜7, graph connectivity and community structure can be analyzed using three types of pivot behavior. In the following, we illustrate how these quantities can be employed to interpret our two-dimensional representations of the graph datasets.
@@
In Figure˜1(b) of the MUTAG dataset, class B has a higher growth rate and smaller pivot than the red class. This shows that the class is growing faster than class A with respect to the closeness function in the MUTAG dataset, i.e., the graph has a low diameter. Similarly, in contrast, in the PROTEINS dataset (Fig. 4(a)), the growth rates are similar for both classes ($\sim 1.5-1.7$), but the pivot (initial graph size) is smaller in class A. This implies that class A has fewer edges in relation to the number of nodes. Such patterns, as described in Sec. B.4, can reveal key insights into graph topology. In a similar vein, TopER visualizations can be used for anomaly detection. For example, in Fig. 1(a), an outlier PROTEINS graph alone has a positive pivot and appears as the rightmost data point.
@@
More importantly, TopER homogenizes graph representations, allowing us to compare graphs across datasets, which may open new paths in training graph foundation models. To our knowledge, TopER is unique in directly producing interpretable 2D embeddings for cross-dataset visualization without relying on learned high-dimensional encodings or opaque projections, unlike GPSE (CLL+, 24) and GFSE (CZW+, 25), which rely on learned high-dimensional embeddings. For example, Figure˜1(a) visualizes graphs of three datasets on the same panel, where we see that Mutag and COX2 differ in their pivot only. The similarity is not surprising; MUTAG and COX2 are datasets of molecular graphs where nodes are atoms and edges are chemical bonds. As the molecules in both datasets have similar types of atoms and bond configurations (e.g., ring structures), TopER captures these similarities, leading to similar embeddings.
These examples highlight that, in many practical settings, TopER’s interpretability outweighs modest performance differences compared to more expressive or data-driven models. As shown in Figure˜4, the pivot–growth representation captures fine-grained structural variations, such as changes in edge density, community organization, and filtration behavior, through simple geometric patterns that can be directly visualized and interpreted. This enables users to pinpoint the topological mechanisms responsible for observed differences between classes or datasets. Beyond classification, such interpretability makes TopER particularly well-suited for applications where structural understanding is critical. Its training-free and computationally efficient design further allows deployment in large-scale or data-scarce environments, as well as integration into hybrid pipelines where TopER embeddings serve as interpretable anchors for downstream learning models. Thus, while deep or spectral approaches may achieve marginally higher benchmark performance in some cases, TopER offers a complementary framework that prioritizes clarity, scalability, and theoretical grounding in topological structure.
Limitations.
While our approach is designed to be efficient and broadly applicable, its performance can vary depending on the choice of filtration function, which may require domain knowledge in certain applications. In practice, we found the method to be robust across a variety of datasets, and further refinements to filtration strategies could enhance adaptability to new domains.